Friday 5th December 2014
Thursday 4th December 2014 at HRC: samples from Mission #30 to 33 (from October 1 to
November 18, 2014) and Forerunner CTD cruise (November 20, 2014): Nutrient samples
deposited into Needoba/Peterson -20 freezer; measured “Raw Fluorescence” (RFU) of
Chlorophyll a samples with Needoba/Peterson Turner Trilogy fluorometer; worked with Tawnya
and Michelle Maier to generate Trilogy standard curve using Turner 16.0 ug/L and xxx ug/L
standards; collected data for standard curve using serial dilutions of standards per JG request.
Wednesday 3rd December 2014 at SATURN-03: collected 4L from NS, MW and SB for FLNTU,
NUT and CHL samples; ran aerated DI (temp @ lab = 18°C; temp @03 = 16°C). Spare rim and
tire picked-up at Astoria Ford; mechanic unable to mount tire in spare tire holder: crank
assembly not functioning. SATURN-04: collected 4L from FT and SB for FLNTU, NUT and CHL
samples; ran aerated DI (temp @ lab = 18°C; temp @04 = 15°C); removed AFT-CO2 to return
for calibration and reagents. Michael set-up WetLabs FLNTU SN 801 for sample processing:
discovered this unit was erroneously referenced as SN 802 in narrative for 20Nov14 Forerunner
CTD cruise; FLNTU data saved to Z:>calibration_data>water_samples_03Dec14. NUT and
CHL samples filtered; AFT-CO2 cleaned; surface-water sample bottles prepared pending
sample collection during 04Dec14 Forerunner CTD cruise.
Tuesday 2nd December 2014: filled aerated DI tank in lab (JG emptied and cleaned yesterday)
and started aeration with 2 pumps immediately upon filling at 0930; acid-rinsing and preparing
for water sampling; assisted JG with preparing for chlorophyll sample analyses on Dec 4 with
Tawnya; driver for Michael to SATURN-01.
Monday 1st December 2014: continued processing CTD, SATURN, and Forerunner data.
-----------------------------------------DECEMBER
-----------------------------------------Wednesday 26th November 2014 on Forerunner: after MPW resolved PC and telemetry issues,
ran aerated DI with bucket elevated on main deck (flowrate = 0.8 gpm) and cleaned optics.
Tuesday 25th November 2014 with JG at SATURN-03: ran aerated DI (including scripts).
Monday 24th November 2014, post tutorial with MW: download and learn software to analyze
and compare ADCP, CTD and SATURN station data.
Friday 21st November 2014 with JG at SATURN-03: MW pump pulled and cleaned; flowmeters,
SB reservoir and instruments cleaned; compressor drained (~100 mL).
Thursday 20th November 2014 Forerunner CTD cruise with JG: casts at SATURN-04 outbound
and SATURN-09 outbound (tide still flooding). Problem with CTD software at SATURN-03
outbound, unable to resolve: no additional casts for the remainder of the cruise. Flowthrough
system continued collecting data but also failed at ????time. Collected 1L surface water
samples at SATURN-04 inbound/outbound, SATURN-09 outbound, SATURN-03
inbound/outbound and SATURN-01; filtered for NUT and CHL upon return to lab.
Wednesday 19th November 2014 with JG at SATURN-04: assisted with pulling SB pump for
cleaning and instrument cleaning. Prepared for Forerunner CTD cruise, including prepping
bucket and 1L plastic, amber bottles for surface water sampling at SATURNS-01,-03,-04 and 09.
Tuesday 18th November 2014 with JG at SATURN-03 and SATURN-04: collected NUT and
CHL samples, collected 4L from each depth for FLNTU testing in lab, and ran aerated DI. Used
only HAWP filters and increased filtration volume from 60 to 120 mL for Chlorophyll a (new
protocol for future samples).
-----------------------------------------NOVEMBER
-----------------------------------------Wednesday 22nd October 2014: ran SBE37 SI#1444 and SI#0294 (recovered from Saturn 02 on
28th September) in DI (= “dirty_DI”), data uploaded to Z drive:
field_notes>2014>instrument_upload_2014.
Tuesday 21st October 2014 with JG: measure and begin Saturn 03 SB replacement cable
assembly.
Monday 20th October 2014 with JG at Saturn 03 to address problem with SB temp sensor:
recovered SB cable assembly, pump and temp sensor, returned to shop for troubleshooting and
new cable.
Friday 17th October 2014 with JG at Saturn 03: cleaned SB pump, flow-meters (NS, MW and
SB), drained air compressor (vol ~100mL). At Saturn 04: cleaned SB pump, pH meter, turbidity
and chlorophyll sensors; recollected and filtered samples.
Thursday 16th October 2014 on Forerunner with MW and JG: CTD cruise to Saturn stations 03,
04, 07 and 01; ran aerated DI post-cruise (ADCP272).
Wednesday 15th October 2014 at Saturn 04: collected samples and ran aerated DI. Forerunner
ft: ran aerated DI, cleaned optics, exchanged SBE43 DO sensor #1059 with SBE43 #2279.
Helped with DI tank test for CTD and SBE43 DO sensors.
Tuesday 14th October 2014 at Saturn 03 with Jo to finish NS pump installation and start
temperature sensor data collection. With Jo and Michael: ran 10 minute DI tank CTD test to
prepare for 16 Oct Forerunner cruise: DO values were low; after cleaning and retesting, DO =
97%.
Monday 13th October 2014: added new pump and cable to NS assembly; installed at Saturn 03
but electrical connection is still pending. NS temperature sensor also returned to Saturn 03 and
attached to terminal block; pending initializing software.
Friday 10th October 2014 at Saturn 03 with Jo: pulled MW pump for cleaning; pulled NS pump
to troubleshoot = cable and pump both not functioning; pump, temp sensor and cable assembly
to shop.
Thursday 9th October 2014 at Saturn 04: ran aerated DI; cleaned FT pump then pH, turbidity
and chlorophyll instruments.
Wednesday 8th October 2014 on Forerunner: ran aerated DI (ADCP258), cleaned optics then
debubbled with raw water. SBE37SM-RS232 SN 03707333: ran dirty DI and dirty seawater
(SW).
Friday 3rd October 2014 at Saturn 04: cleaned SB pump, AFT pCO2, CDOM, turbidity and
chlorophyll optics. SBE 37SMP_IDO_R232_03709209 from recovered Saturn 02 buoy:
uploaded data placed in correct folder; ran dirty DI and SW.
Thursday 2nd October 2014: continuing post buoy recovery tasks (cleaning, re-assemble cages,
etc). Forerunner: ran aerated DI (ADCP255), cleaned optics, lowered turbidity sensor in
housing, and de-bubbled system with raw water.
Wednesday 1st October 2014 to Saturn 04: collected samples; ran aerated DI; and cleaned SB
flowmeter.
-----------------------------------------October
-----------------------------------------Tuesday 30th September 2014: uploaded data from SBE37SM-RS232 SN 03707333
recovered from OGI01
(Z>field_notes>2014>instrument_upload_2014>OGI01_data_SBE37SM_0370733>29Sep2014)
then continued instrument and buoy cleaning. Saturn 03: pulled NUT and CHL samples and
ran aerated DI.
Monday 29th September 2014: post-recovery cleaning of Saturn 02 and OGI01
instruments (MW retrieved buoys over the weekend).
Friday 26th September 2014 to Cathlamet Bay with Jo on R/V Corie: for Lydie to collect
sediment samples. Saturn 03 with Jo to clean instruments; pulled MW pump for cleaning.
Thursday 25th September 2014 to Saturn 03 with Jo: to minimize interfering with ESP sampling,
NUT and CHL samples collected from 1140 to 1145 PST for NS, MW and SB and aerated DI
ran between 1240 and 1245 PST. Driver for Michael to Saturn 01.
Wednesday 24th September 2014 to Saturn 04 with Jo: collected NUT and CHL samples for
FT, SB and DI; Jo ran aerated DI and cleaned instruments; pulled FT pump for cleaning and
increased slack in cable assembly to keep pump underwater during extreme low tides.
Tuesday 23rd September 2014: contacted Mary Mosser re fish-finder recovered from Jetty A,
her tech, Steve, will contact us re disposition of this unit and future recoveries and possible
deployments. To Saturn 04 with Jo after power issues were resolved for the SB pump and temp
sensor. Installed terminal block for temp sensor then deployed new cable and both instruments.
Monday 22nd September 2014 on Forerunner: ran aerated DI through ft system (ADCP242),
cleaned optics. lowered turbidity sensor in housing to improve monitoring ran raw water through
system to de-bubble, then another aerated DI run to check turbidity sensor (ADCP243).
Assisted Jo with CTD prep for Forerunner cruise. To Saturn 04 with Jo to install new cable for
SB, deploy refurbished pump and return temperature sensor: power issues with both pump and
temp sensor and troubleshooting efforts were not successful onsite; both instruments returned
to shop.
Friday 19th September 2014 with Jo on Corie to handle boat and assist Michael as he replaced
instruments at Jetty A and recovered NOAA fish-finder; cleaned Corie.
Thursday 18th September 2014 to Saturn 03: collected NUT and CHL samples from NS, MW
and SB with Jo; Jo ran aerated; pulled NS pump for cleaning.
Tuesday 16th September 2014 to Saturn 04: collected NUT and CHL samples from FT and SB;
pulled SB pump for cleaning; SB cable returned to shop to troubleshoot temperature sensor
failure. To Saturn 01 with Michael.
Monday 15th September 2014 on Forerunner to run aerated DI (ADCP230) and clean
instruments with Jo.
Friday September 12th 2014 at Saturn 03 with Jo to troubleshoot problem with MW and SB flow:
air compressor had failed, returned to shop. Boat-handler for Michael to Saturn 09: replaced
curly cable with heavier gauge cable and battery pack.
Thursday September 11th 2014 at Saturn 03 with Jo: cleaned MW pump; replaced SB pump;
cleaned SB overflow reservoir and pCO2 AFT (returned to lab: out of reagent and detector cage
is damaged).
Wednesday September 10th 2014 on Forerunner: ran aerated DI (ADCP227) and cleaned
instruments with Jo. Problem with MW flow at Saturn 03: cleaned flowmeter to resolve. Boathandler for Michael to Saturn 09: replaced solar charge controller.
Tuesday September 9th 2014 with JG at Saturn 03: collected samples while Jo ran aerated DI.
Boat-handler for Michael to Saturn 09 installed solar panel, cable and telemetry.
Monday September 8th 2014 at Saturn 04 to lower FT pump cable: gaps in data due to pump
being out of water at low tides. Boat-handler to Saturn 09 with Michael while he installed
instruments and float frame.
Friday September 5th 2014: updated inventories, sample logsheets, familiarized with data
outputs from instruments, and re-processed Forerunner flowthrough system DI data.
Thursday September 4th 2014: boat handling session on Corie with Ethan (east basin docking,
crossing channel). Saturn 04: collected NUT and CHL samples (began using silicone gaskets
in Swinnex holders when collecting CHL samples with HAWP filters); ran aerated DI; cleaned
FT pump.
Wednesday September 3rd 2014: driver for Michael to Saturn01. Boat-handling session on R/V
Corie with Ethan (east basin to Saturn 04). Corie boat-handler for Michael and Ethan to
continue installation of station infrastructure at Saturn 09 (cables and lump).
Tuesday September 2nd 2014: worked on Forerunner ADCP flowthrough system DI data
-----------------------------------------September
-----------------------------------------Friday September 29th 2014 on Forerunner: ran aerated DI (ADCP214), retrieved data for
processing, cleaned PE, CDOM, Chlorophyll and Turbidity sensors; determined Turbidity sensor
info: Turner CYCLOPS-7T SN 2101232; reconciled instruments with circuit board. Saturn 03 to
troubleshoot SB pump flow: no discharge from SB overflow reservoir, pulled pump and
discovered it was cycling, cleaned dirty flowmeter (noticed it had been stuck), redeployed pump,
observed flowmeter was spinning at slow rate with little overflow from reservoir: flow from pump
is still feeble.
Thursday August 28th 2014 at Saturn 04: sampled and ran aerated DI (only one successful
SUNA sample). Noticed AFT pCO2 inlet/outlet tubing is reversed: will correct during next
cleaning.
Wednesday August 27th 2014 with JG at Saturn 03: sampled NS, SB and DI; MW pump not
working, after JG confirmed no problems with electrical and we observed cable was in good
condition, pump was pulled and replaced with refurbished pump from lab. Noticed the
replacement pump was not working well. Pulled SB pump to troubleshoot problem with
decreased flowrate: pump was operational, so removed barnacles from intake screen, resulted
in low flow with slow-moving flowmeter; will continue to ascertain cause of inconsistent flow.
Tuesday August 26th 2014 with MW and JG on R/V Corie: TRIED to assist with first phase of
Saturn 09 infrastructure; sheathed/taped exposed waterline tubing on walkway at Saturn 03.
Monday August 25th 2014 per Field Group Meeting: tasked with Forerunner DI data, instrument
inventory, and NUT/CHL sample logsheet.
Thursday August 21st 2014: drove MW to Saturn01 to install reworked conductivity cell cable.
Wednesday August 20th 2014: uploaded Saturn 03 ADP biofouling images (retreived August 5th
2014): CCALMR19 >data> digital images>Saturn-03>ADP_frame biofouling. Began
reorganizing MERTS lab with the addition of 4 bookcases (via MW).
Tuesday August 19th 2014 with Jo at Saturn 04 collected and filtered NUT and CHL samples ran
aerated DI; assisted Jo with removal of pH probe and controller (to shop for calibration).
Learned to calibrate pH instruments with Jo then returned probes and controllers to Saturn 03
and 04.
Monday August 18th 2014 with JG at SAT-03: collected and filtered NUT and CHL samples; Jo
ran aerated DI; assisted Jo with removal of pH probe and controller (to shop for calibration).
On Forerunner with MW and JG: inventoried instruments MW installed for flowthrough (ft)
system: thermosalinograph SeaBird 45 SN 4551581-0252, Phycoerythrin fluorometer Turner
CYCLOPS-7 E SN 2180540, CDOM fluorometer Turner CYCLOPS-7 U SN 2180542,
Chlorophyll fluorometer Turner CYCLOPS-7C SN 2101235, Turbidity sensor (TBD), and
Dissolved Oxygen sensor Seabird 43 SN 431059; MW ran aerated DI (ADCP209) and uploaded
data (from thumbdrive).
Thursday August 14th 2014 with JG at Saturn 04: collected and filtered NUT and CHL samples
(note: CHL samples = 1 ea GF/F and 1 ea HAWP at each depth); ran aerated DI; cleaned SB
pump; cleaned Turner Cyclops turbidity sensor and chlorophyll fluorometer.
Wednesday August 13th 2014 with JG at Saturn 03: collected and filtered NUT and CHL
samples (note for today only: CHL samples = 2 ea GF/F and 2 ea HAWP filters at ea depth plus
DI = 16 filters total).
Friday August 8th 2014 with MW and JG on Corie: assist MW while measuring for Saturn 09
configuration under Young’s Bay Bridge; helped Jo replace MW pump cable at Saturn 03.
Thursday August 7th 2014 with JG at Saturn 03: PC issues precluded cleaning NS pump;
cleaned reservoir for SB pump overflow while Jo cleaned turbidity sensor; cleaned FT pump and
pH probe at Saturn 04.
Wednesday August 6th 2014 with JG: collected NUT and CHL samples at Saturn 03 and Saturn
04. GF/F filters continue to be used for NUT sample filtrate. GF/F and 0.2um HAWP (no orings) filters used for CHL samples.
Tuesday August 5th 2014 on Corie with JG and MW and with Forerunner assistance: support
recovery and deployment of ADP at Saturn03. Photodoc biofouling of retreived ADP.
Friday August 1st 2014 on Corie (with MW and JG): removed extraneous ropes and bungees
from Sat03 piling pending ADP deployment; Sat07 for Michael to resurrect CT pump (clear air in
pump) and replace antenna; Sand Island for Michael to troubleshoot (solar controller needs to
be replaced); Jetty A for Michael to install relay.
-----------------------------------------AUGUST
-----------------------------------------Thursday 31 July 2014 at Sat03 (with MW and JG): retrieve ESP for return to MERTS and
Michael to troubleshoot MW pump (cable flooded); ESP pucks to Lydie.
Wednesday 30 July 2014 on Forerunner (with Vanessa and interns) to Sat07: retrieve buoy for
Michael to replace CT/clean instruments then redeployed (solar panel glass and antenna
damaged). On Corie (with MW and JG) to Sat03: aborted retrieving old/deploying new ADP
(Foreunner Captain concerned about conditions); to John’s Bay Bridge: photodocument
substructure for Michael to submit Sat09 instrument permit diagram.
-----------------------------------------JULY
------------------------------------------
Tuesday 17 June 2014: Sat04: collected nutrient and chlorophyll samples from FT and SB; ran
aerated DI (need to request root access).
Monday 16 June 2014 with JG: Sat03: collected NUT and CHL samples; problem with aerated
DI run.
Friday June 13 2014 with JG: Sat03: ran aerated DI; cleaned MW pump with Triton X and
bleach; cleaned instruments (ISUS 153, CDOM 1337, Algaewatch, and cyanowatch); cleaned
reservoir for SB pump. SAT04: cleaned instruments (Cyclops TB 2180-448, Cyclops C 2100956, and TSG 0109).
Thursday June 12 2014 with JG and MW on Forerunner: deployed SBE SMP #7761 at CBNC3.
Sat03 with JG: collected nutrient (NUT) and chlorophyll (CHL) samples; script issues prevented
running aerated DI; cleaned AFT pCO2. Sat04 with JG: collected NUT and CHL samples; ran
aerated DI; cleaned AFT pCO2.
Friday 6 June 2014 with JG: Sat03: collected nutrient and chlorophyll samples; ran aerated DI;
further work thwarted by overpowering odor of ammonia and increased activity of cannery
workers. Sat04: collected nutrient and chlorophyll samples; ran aerated DI; reinstalled AFT
pCO2 C-0007; cleaned CDOM.
Thursday June 5 2014 with JG and MW: videoconferenced NSF site visit
Wednesday June 4 2014 with JG: Michael to/from Sat01; to Long Beach to collect
lithium battery order.
-----------------------------------------JUNE
-----------------------------------------Friday 30 May 2014 with JG: Sat03: collected nutrient and chlorophyll samples from
NS, MW, and SB pumps and DI; ran aerated DI. Sat04: collected nutrient and
chlorophyll samples from FT and SB pumps; ran aerated DI; cleaned X,X,X and X.
Thursday 29 May 2014: Forerunner cruise: CTDs Sat01, Sat03, Sat04 and Sat07.
Buoy at Sat07 brought on deck for cleaning and buoy recovery from Cathlamet Bay.
Wednesday 28 May 2014: with JG prepped for Forerunner cruise; tested CTD in
aerated DI tank; attempted to prep SMP XXXX for deployment at Cathlamet Bay
(CBNC3) but wire disconnected from XXXX.
Friday 23 May 2014: set-up AFT-CO2 with DI water to run over the weekend.
Thursday 22 May 2014: Sat03 with JG: collected nutrient and chlorophyll samples from
seabed, mid-water and near-surface pumps plus DI; ran aerated DI; replaced nearsurface pump before it was discovered that midwater pump was intended for
replacement: resolved by replacing mid-water pump with the just-removed near surface
pump. Sat04 with JG: collected nutrient and chlorophyll samples from seabed and float
pumps; ran aerated DI; cleaned seabed pump with TritonX and bleach; cleaned pH
probe; removed AFT-CO2 SN AC0007 for maintenance at MERTS (cleaned, installed
new reagents, and downloaded data).
Friday 09 May 2014: training with MW re deployment of Wendy Smythe's CT in Soda
Bay. Aggregated and listed nutrient samples missions #1 to 9 for analysis in Portland;
transported on ice and stored in Needoba -20 in Hatfield Research Center (HRC3).
Thursday 08 May 2014: Sat04 with JG: collected nutrient and chlorophyll samples from
seabed and float pumps and DI; ran aerated DI, pulled and cleaned float pump with
TritonX and bleach
Wednesday 07 May 2014: to Sat03 with MW/KR/JG to retrieve ESP, sample collection
and maintenance (run aerated DI), but Point Adams Cannery activities prevented
access to the station. Aborted.
Thursday 01-May-3014: Sat04 with JG: collected nutrient and chlorophyll samples
from seabed and float pumps, ran aerated DI (run interrupted, needs to be redone),
pulled and cleaned seabed pump with TritonX and bleach
--------------------------
MAY
--------------------------
Wednesday 30-Apr-2014: Sat03 with JG: collected nutrient and chlorophyll samples at
all 3 depths, ran aerated DI; observed JG removing AFT-CO2 for cleaning and reagent
replacement in lab
Thursday 24-Apr-2014: Sat04 training with JG: collected nutrient and chlorophyll
samples at both depths, ran aerated DI, cleaned all instruments
Wednesday 23-Apr-2014: Sat03 training with JG: collected nutrient and chlorophyll
samples at all depths, ran aerated DI
Friday 18-Apr-2014: helped Lydie prep ESP for deployment
Thursday 17-Apr-2014: observed JG : Sat03 run aerated DI and clean instruments;
Sat04 clean FT pump with TritonX and bleach, run aerated DI
Wednesday 16-Apr-2014: Sat04 observed JG install pH probe and analyzer, and
CDOM
Thursday 11-Apr-2014: Forerunner cruise with MW: CTDs at Saturn 01, 03, 04, 07;
retrieved buoy at Saturn07
Wednesday 10-Apr-2014: observed MW preparing CTD for Forerunner cruise; with MW
at Sat03 to troubleshoot
Download

sheehan_field_notes_2014